OpenStudy (anonymous):

Under what conditions does sqrt{x ^{2}+ y ^{2} = x + y ? Please explain why, also.

OpenStudy (anonymous):

\[\sqrt{x ^{2}+y ^{2}} = x + y\]

OpenStudy (zarkon):

look at \[x^2+y^2=(x+y)^2\]